# Multi-temporal crack segmentation in concrete structure using deep learning approaches

### v1

@@ 4.2.3 Evaluation of false positives:
The segmentation masks clearly show that Swin UNETR is more effective at identifying thin cracks than U-Net. However, it is important to note that the performance of both models for thin and low-contrast cracks is lower than that for more common cracks.
@@ 5 Conclusions
This study examined the effect of multi-temporal data on the performance of semantic segmentation for crack images. The Swin UNETR was trained on a multi-temporal crack dataset, whereas a U-Net was trained on the deserialized version of the same dataset for comparison. Swin UNETR achieved an Intersection over Union (IoU) of $82.72\%$ and an F1-score of $90.52\%$, whereas U-Net resulted in an IoU of $76.69\%$ and an F1-score of $86.81\%$ on the test set. This indicates that the multi-temporal approach outperformed the mono-temporal approach, even when utilizing only half the parameters compared to U-Net.
The multi-temporal Swin UNETR demonstrated superior segmentation quality compared with the mono-temporal U-Net. Swin UNETR predictions exhibited greater consistency, reduced noise, and enhanced ability to distinguish cracks from visually similar features in the images. By leveraging the entire time-series data, the multi-temporal model significantly improved its performance in ambiguous image regions, resulting in more confident predictions. In contrast, mono-temporal U-Net encountered several challenges. It struggled to segment larger cracks without introducing artifacts, and often produced noisy predictions. Furthermore, U-Net frequently failed to differentiate between actual cracks and visually similar features, highlighting its limitations in this specific application. Access to temporal information proved to be a crucial advantage for the Swin UNETR, enabling it to make informed decisions based on the evolution of features over time. This temporal context allowed the model to better interpret ambiguous areas and reduce false positives, resulting in more accurate and reliable crack segmentation.
Acknowledgment: The authors would like to thank the DFG (German Research Foundation) for the support within the project “Optical 3D-Bridge-Inspect” – Project Number 501682769 as part of DFG’s Priority Programme 2388 “Hundred plus”.

### v2

@@ 4.2.3 Evaluation of false positives and false negatives:
False negatives in this context are cracks which were not identified by the models. The Swin UNETR has a higher recall than the U-Net and therefore is able to find more cracks, like in Figure 6 d). However, it is important to note that the performance of both models for thin and low-contrast cracks is lower than that for more common cracks.
@@ 5 Conclusions
This study examined the effect of multi-temporal data on the performance of semantic segmentation for crack images. The Swin UNETR was trained on a multi-temporal crack dataset, whereas a U-Net was trained on the deserialized version of the same dataset for comparison. Swin UNETR achieved an IoU of $82.72\%$ and an F1-score of $90.52\%$, whereas U-Net resulted in an IoU of $76.69\%$ and an F1-score of $86.81\%$ on the test set. This indicates that the multi-temporal approach outperformed the mono-temporal approach, even when utilizing only half the parameters compared to U-Net.
The multi-temporal Swin UNETR demonstrated superior segmentation quality compared with the mono-temporal U-Net. Swin UNETR predictions exhibited greater consistency, reduced noise, and enhanced ability to distinguish cracks from visually similar features in the images. By leveraging the entire time-series data, the multi-temporal model significantly improved its performance in ambiguous image regions, resulting in more confident predictions. These properties enable a use-case of multi-temporal segmentation models in SHM in long-term scenarios. On one hand such models have less false negatives, thereby triggering fewer false alarms and on the other hand the superior segmentation quality facilitates a better detection of actual cracks. Due to available temporal information, small and continuous misalignments of images will be learned by the model, making it easier for deployment in real-life scenarios. By employing data augmentation during training, models can be adapted to be robust against changes to the surface due to weathering.
In contrast, mono-temporal U-Net encountered several challenges. It struggled to segment larger cracks without introducing artifacts, and often produced noisy predictions. Furthermore, U-Net frequently failed to differentiate between actual cracks and visually similar features, highlighting its limitations in this specific application. Access to temporal information proved to be a crucial advantage for the Swin UNETR, enabling it to make informed decisions based on the evolution of features over time. This temporal context allowed the model to better interpret ambiguous areas and reduce false positives, resulting in more accurate and reliable crack segmentation.
Acknowledgment: The authors would like to thank the DFG (German Research Foundation) for the support within the project “Optical 3D-Bridge-Inspect” – Project Number 501682769 as part of DFG’s Priority Programme 2388 “Hundred plus”. This research was supported by DFG (German Science Foundation) through a Major Research Instrumentation, project number 461109100.